# SpecGen: Automated Generation of Formal Program Specifications via Large Language Models

### v2

@@ V-C RQ3: Effectiveness of Selection Strategy
Performance on SpecGenBench. Using the random selection strategy takes 36.20 verifier calls on average while the heuristic selection strategy takes 28.51 calls in terms of five categories in SpecGenBench. Hence, the heuristic selection strategy achieves an improvement of 21.23%. Furthermore, we can observe that the improvements in different categories of programs vary significantly. The improvements in the loop categories including single-path, multi-path, and nested loop are more significant than sequential and branched. The main reason is that generating specifications for loop-containing programs is challenging, and usually requires more iterations to obtain verifiable specifications. In this case, a good selection strategy often highlights advantages more effectively. However, the improvements in sequential and branched categories are fewer. The reason is similar to SV-COMP. The high efficiency of conversational generation on sequential and branched programs makes the selection strategies invalid. Nevertheless, loop structures are common in programs, thus a heuristic selection strategy to improve the validation efficiency is still helpful and necessary.
RQ3: The heuristic selection strategy effectively improves the efficiency of SpecGen. It is especially useful to generate the specifications for programs that have more complex structures such as loops.
@@ V-D RQ4: User Study on the Quality of Specifications
To evaluate the semantic quality of the generated specifications, we conduct a user study for evaluation. 15 Ph.D. students are invited to rate the specifications generated by different approaches. A detailed description of the questionnaire and rating criteria is given in Section IV-D. We selected the 15 programs from the dataset of SpecGenBench that can be handled by all of Houdini, Daikon, and SpecGen. Apart from the specifications generated by these approaches, we also add the ground truth as a reference. The rating scores are presented in Table V, where the average score for full marks is 5.
We can observe that the ground truth specifications (oracle) receive an average rating score of 4.83, indicating that the semantics of these programs can be described comprehensively through JML specifications. Furthermore, the specifications generated by SpecGen received a rating score of 4.54, which is close to the oracle, indicating that the generated specifications by SpecGen can also describe the real behaviors of the input program more fully. Among the 15 programs, all rating scores given to SpecGen are above 3, with the lowest rating being 3.33, meaning that in the worst case, SpecGen can still generate non-trivial specifications about the properties of the input program. Benefiting from the code comprehension capability of LLMs, SpecGen can generate specifications closely related to the real behaviors of the input program. In comparison, the specifications generated by Houdini and Daikon received an average rating score of 2.32, reflecting the semantic weakness in these specifications, which is consistent with the claims in Section II-B. As Houdini and Daikon rely on pre-defined templates for generation, the specifications produced by them are often simplistic and trivial, making it challenging to precisely capture the actual behavior and functionality of the input program. We further provide a rating of 100 programs that SpecGen can successfully generate verifiable specifications following the criteria from Section IV-D on our website [17].
@@
VI Threats to Validity
@@ VII Related Work
Program Specification Generation. The research on program specification generation can be categorized into two types: natural language specification generation and formal specification generation. Natural language specification generation primarily manifests as code summarization [53], a process of automatically generating accurate, human-readable descriptions of code functionality. Numerous efforts have been made to utilize machine learning methods for code summarization [54, 55, 56, 57]. Formal specification generation primarily takes the form of the generation of program invariants, the formal language representations of properties that a program is guaranteed to satisfy at a certain program point. In invariant generation, a large amount of research focuses on the generation of loop invariants [58, 59, 60, 61, 62, 63, 64], while the rest of the works attempt to generate invariants of other forms, e.g. pre-conditions [65, 42], post-conditions [66, 67, 36, 68, 69], assertion-based invariants [70] and finite automata [71, 72, 73]. There also have been efforts employing fine-tuning methods to help LLMs better adapt to specification generation [74]. Among these works, artifacts of Ghosal et al. [42] and Pei et al. [74] are not publicly available, the artifact of Alshnakat et al. [36] is built for C code and Frama-C contracts, and the grammar for specifications of Molina et al. [68, 69] involves specific features, which cannot be trivially translated into equivalent JML, thus we cannot include them for comparison. SpecGen utilizes the code comprehension capability of LLMs for program specification generation in a conversational manner, further followed by the mutation-based approach for enhancement.

### v3

@@ V-C RQ3: Effectiveness of Selection Strategy
Performance on SpecGenBench. Using the random selection strategy takes 36.20 verifier calls on average while the heuristic selection strategy takes 28.51 calls in terms of five categories in SpecGenBench. Hence, the heuristic selection strategy achieves an improvement of 21.23%. Furthermore, we can observe that the improvements in different categories of programs vary significantly. The improvements in the loop categories including single-path, multi-path, and nested loop are more significant than sequential and branched. The main reason is that generating specifications for loop-containing programs is challenging, and usually requires more iterations to obtain verifiable specifications. In this case, a good selection strategy often highlights advantages more effectively. However, the improvements in sequential and branched categories are fewer. The reason is similar to SV-COMP, where the high efficiency of conversational generation on sequential and branched programs makes the selection strategies invalid. Nevertheless, loop structures are common in programs, thus a heuristic selection strategy to improve the validation efficiency is still helpful and necessary.
RQ3: The heuristic selection strategy effectively improves the efficiency of SpecGen. It is especially useful when generating specifications for programs with more complex structures such as loops.
@@ V-D RQ4: User Study on the Quality of Specifications
A user study is conducted to evaluate the semantic quality of the generated specifications. 15 Ph.D. students are invited to rate the specifications generated by different approaches. A detailed description of the questionnaire and rating criteria is given in Section IV-D. We selected the 15 programs from the dataset of SpecGenBench that can be handled by all of Houdini, Daikon, and SpecGen. Apart from the specifications generated by these approaches, we also add the ground truth as a reference. The specifications are kept anonymous to the students, disclosing no information about the sources of the specifications. The rating scores are presented in Table V, where the average score for full marks is 5.
We can observe that the ground truth specifications (oracle) receive an average rating score of 4.83, indicating that the semantics of these programs can be described comprehensively through JML specifications. Furthermore, the specifications generated by SpecGen received a rating score of 4.54, which is close to the oracle, indicating that the generated specifications by SpecGen can also describe the real behaviors of the input program more fully. Among the 15 programs, all rating scores given to SpecGen are above 3, with the lowest rating being 3.33, meaning that in the worst case, SpecGen can still generate non-trivial specifications about the properties of the input program. In comparison, the specifications generated by Houdini and Daikon received an average rating score of 2.32, reflecting the semantic weakness in these specifications. Houdini and Daikon rely on pre-defined templates, which are in fact independent from the input program and can only cover a limited number of specification patterns. Consequently, the specifications produced are often simplistic and trivial, involving only a narrow range of variables and operators, making it difficult to capture the actual behavior and functionality of the input program precisely. Unlike traditional approaches that rely on a fixed set of templates, SpecGen utilizes the code comprehension capabilities of LLMs, which can cover a larger range of scenarios and generate targeted specifications that more closely match the semantics of the input program. We further provide a rating of 100 programs that SpecGen can successfully generate verifiable specifications following the criteria from Section IV-D on our website [22].
@@
VI Discussion
VI-A Performance on Real-world Programs
TABLE VI: Performance on programs collected from 9 repositories in Defects4J.
Approaches
chart
cli
codec
compress
jackson
jxpath
lang
math
time
Total
Num. Prob. Num. Prob. Num. Prob. Num. Prob. Num. Prob. Num. Prob. Num. Prob. Num. Prob. Num. Prob. Num. Prob.
Daikon 3 - 3 - 0 - 1 - 3 - 2 - 2 - 1 - 0 - 15 -
4-shot LLM 1 7.14% 2 7.33% 2 9.71% 3 11.67% 3 9.52% 1 2.78% 4 11.67% 3 17.50% 1 5.00% 20 8.97%
Conversational 4 47.62% 4 60.00% 3 41.67% 1 11.11% 2 19.05% 5 55.56% 3 28.57% 3 66.67% 3 41.67% 28 39.33%
SpecGen 6 68.57% 4 68.00% 4 65.00% 4 36.67% 4 42.86% 5 63.33% 5 65.71% 3 45.00% 3 35.00% 38 55.20%
To further evaluate the performance of SpecGen on real-world programs, we collect programs involved in Defects4J [50], a well-known dataset of reproducible bugs within open-source repositories. During the collection process, we only consider individual files with no dependency on third-party libraries or other files in the repository. This ensures that all the collected files can be properly executed and verified outside the repository. Eventually, 50 Java source files from 9 repositories are collected. The average line of code and cyclomatic complexity of the collected programs are 374.78 and 18.29, respectively. We follow the same experimental settings in Section IV-A for experiments. Note that we only aim to evaluate the verifiability of the generated specifications in the same way Section V-A does, so the ground truth specifications of the programs are not prepared.
Table VI shows the performance of SpecGen and other baseline methods on the programs extracted from Defects4J. Although Daikon underperforms the LLM-based approaches, it still exhibits certain abilities in processing real-world programs. This is due to the existence of some simplistic methods within real-world class definitions, such as those retrieving the value of a certain class member without doing anything else, which Daikon is capable of handling. Compared to Daikon, the LLM-based approach with the simplest setting, i.e. 4-shot, achieved 5 more programs handled. Based on the few-shot learning technique, the conversational approach further achieved 28 programs handled. Lastly, SpecGen succeeds in handling 38 out of the 50 programs, with an average success probability of 55.20%, displaying decent capabilities in handling real-world programs.
VI-B Threats to Validity
@@ VII Related Work
Program Specification Generation. The research on program specification generation can be categorized into two types: natural language specification generation and formal specification generation. Natural language specification generation primarily manifests as code summarization [60], a process of automatically generating accurate, human-readable descriptions of code functionality. Numerous efforts have been made to utilize machine learning methods for code summarization [61, 62, 63, 64]. Formal specification generation primarily takes the form of the generation of program invariants, the formal language representations of properties that a program is guaranteed to satisfy at a certain program point. In invariant generation, a large amount of research focuses on the generation of loop invariants [65, 66, 67, 68, 69, 70, 71], while the rest of the works attempt to generate invariants of other forms, e.g. pre-conditions [72, 48], post-conditions [73, 13, 41, 74, 10], assertion-based invariants [75] and finite automata [76, 77, 78]. With the development of Large Language Models, there have also been efforts employing LLMs to generate program specifications. Wen et al. [21] combine LLMs with static analysis techniques, including code decomposition, to generate verifiable program specifications. Pei et al. [79] utilize fine-tuning to enhance the performance of LLMs on specification generation tasks. Concerning the difficulty of selecting correct specifications from the massive LLM-generated results, Chakraborty et al. [69] propose a ranking algorithm that can distinguish correct inductive invariants from incorrect attempts based on the problem definition. Among these works, artifacts of Ghosal et al. [48] and Pei et al. [79] are not publicly available, the artifact of Alshnakat et al. [41] is built for C code and Frama-C contracts, and the grammar for specifications of Molina et al. [74, 10] involves specific features, which cannot be trivially translated into equivalent JML, thus we cannot include them for comparison. Compared to these works, SpecGen utilizes the code comprehension capability of LLMs for program specification generation in a conversational manner, further followed by the mutation-based approach for enhancement.